A. Jensen Tire & Auto is in the process of deciding whether to purchase a maintenance contract for its new computer wheel alignment and balancing machine. Managers feel that maintenance expense should be related to usage, and they collected the following information on weekly usage (hours) and annual maintenance expense (in hundreds of dollars): Weekly Usage (hours) Annual Maintenance Expense ($100) 13 17.0 20 30.0 28 37.0 32 47.0 17 30.5 24 32.5 40 51.5 38 40.0 A3. Using the OLS results and find the residual, i.e., ei, that corresponds with 32 hours of weekly usage of this machine: a. 37.2004 b. 42.2348 c. 5.7652 d. 2.1963
Answer #1

y: Annual Maintenance Expense ($100)

x: Weekly Usage (hours)

=a +bx

' 'a' indicates the value of Y when X=0. It is called Y-intercept

'b' indicates the slope of the regression line and gives a measure of change of 'Y' for a unit change in 'X'

x: WeeklyUsage(hours) y: AnnualMaintenanceExpense xy x2
13 17.0 221 169
20 30.0 600 400
28 37.0 1036 784
32 47.0 1504 1024
17 30.5 518.5 289
24 32.5 780 576
40 51.5 2060 1600
38 40.0 1520 1444
=212 =285.5 =8239.5 =6286

= 8.959394 + 1.008608x

Residual : Actual y - predicted y: y-

x:WeeklyUsage(hours) y:AnnualMaintenanceExpense =8.959394+1.008608x Residual=y-
13 17.0 22.071298 -5.071298
20 30.0 29.131554 0.868446
28 37.0 37.200418 -0.200418
32 47.0 41.23485 5.76515
17 30.5 26.10573 4.39427
24 32.5 33.165986 -0.665986
40 51.5 49.303714 2.196286
38 40.0 47.286498 -7.286498

Residual for 32 = 5.765155.7652

residual, i.e., that corresponds with 32 hours of weekly usage of this machine:

c. 5.7652